Assume that Janet Leigh purchased stock that costs $12.25 per share and purchased 1,000 shares; If...

  1. Assume that Janet Leigh purchased stock that costs $12.25 per share and purchased 1,000 shares;

  1. If Janet sold all shares for $21.75 after holding the stock for 15 months, and the single taxpayer was in the 39.6% bracket what are the taxes on this gain?

  1. If Janet sold all shares for $19.50 after holding the stock for 8 months, and the single taxpayer was in the 39.6% bracket, what are the taxes on this gain?
Answer #1

PURCAHSE COST FOR SHARE= 1000*12.25 = 12250.

SOLD FOR 21.75.

A. 1000*21.75=21750.

PROFIT: 21750-12250=9500.

TAX ON GAIN : 9500*39.6%=3762.

B. 1000* 19.50= 19500.

PROFIT: 21750-19500=2250.

TAX ON GAIN: 2250*39.6%=891.